OpenStudy (anonymous):

Calculate the integral, if it converges. You may calculate the limit by appealing to the dominance of one function over another, or by l'Hopital's rule. The integral is sinx/sqrt(cosx) from pi/4 to pi/2. How do i solve this?

OpenStudy (anonymous):

\[\int\limits_{\pi/4}^{\pi/2} \sin(x)/\sqrt \cos(x)\]

OpenStudy (anonymous):

since cos(pi/2) = 0, we have \[\lim t \rightarrow \pi/2 \int\limits_{\pi/4}^{t} \sin(x)/\sqrt{\cos(x)} dx\]

OpenStudy (anonymous):

to do the integral, set u = cos(x), du = -sin(x) dx

OpenStudy (anonymous):

can take it from here?

OpenStudy (anonymous):

yep, i got it! \[2\sqrt(\sqrt(2)/2\] thanks!